# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 9.2 $G=SL_{2n},H=Sp_{2n}$
Let $G=SL_{n}(F)$, $H=Sp_{2n}(F)$ is a symmetric subgroup of $G$. $X=G/H$ can be identified with anti symmetric matrices in $G$.
Let $T\subset G$ be the maximal torus of diagonal matrices. Embed $W_{aff}=N_{G}(T)/T^{0}$ into $G$ as the group of permutation matrices whose non zero entries are powers of the uniformazer $\pi$ (up to a sign).
@@
  • In this case there is a unique orbit for the action of $H$ on the $\sigma$ stable apartments in $\mathcal{B}_{G}$. The $I$ orbits on $X$ corresponds to anti symmetric matrices in $W_{aff}$.
  • The action of $W_{aff}$ on $I\backslash X$ is given by $w\times x=wxw^{t}$.
@@
    $T_{s}1_{x}=1_{sxs^{t}}\text{ if }l(sxs^{t})>l(x)$
    $T_{s}1_{x}=q1_{sxs^{t}}+(q-1)1_{x}\text{ if }l(sxs^{t})<l(x)$
    $T_{s}1_{x}=q1_{x}\text{ if }sxs^{t}=x$
  • In the case $n=2$, $W_{aff}$ is generated by four simple reflections $s_{0},s_{1},s_{2}$ and $s_{3}$. The hyper graph $\Gamma_{X}$ is the graph appearing in figure 2 continued downward periodically.
    Figure 2: Part of the hyper graph $\Gamma_{X}$ of the symmetric space $SL_{4}/Sp_{4}$. Each hyper edge is either of size 2 or 1.

### v3

@@ 9.2 $G=SL_{2n},H=Sp_{2n}$
[…]
Let $T\subset G$ be the maximal torus of diagonal matrices. Embed $W_{aff}=N_{G}(T)/T^{0}$ into $G$ as the group of permutation matrices whose non zero entries are powers of the uniformizer $\pi$ (up to a sign).
@@
  • There is a unique orbit for the action of $H$ on the $\sigma$ stable apartments in $\mathcal{B}_{G}$. The $I$ orbits on $X$ correspond to antisymmetric matrices whose non zero entries above the diagonal are powers of the uniformizer $\pi$. i.e. we can identify $X$ with the space of antisymmetric matrices and every $I$ orbit on $X$ contains a unique element of the mentioned form.
    Denote by $nor:AntiSymmetric\rightarrow X/I$ the map that takes an antisymmetric matrix and replaces every entry $a$ above the diagonal with $\pi^{\nu(a)}$ and every entry $a$ below the diagonal by $\pi^{\nu(a)}$. Here, $\nu$ is the valuation of $F$.
  • The right action of $W_{aff}$ on $X/I$ is given by $x\times w=nor(w^{t}xw)$.
@@
    $1_{x}T_{s}=1_{s^{t}xs}\text{ if }l(s^{t}xs)>l(x)$
    $1_{x}T_{s}=q1_{s^{t}xs}+(q-1)1_{x}\text{ if }l(s^{t}xs)<l(x)$
    $1_{x}T_{s}=q1_{x}\text{ if }s^{t}xs=x$
  • In the case $n=2$, $W_{aff}$ is generated by four simple reflections $s_{0},s_{1},s_{2}$ and $s_{3}$. The hyper-graph $\Gamma_{X}$ is the graph appearing in Figure 2 continued downward periodically.
[…]